# Uncertainty-Aware Search and Value Models: Mitigating Search Scaling Flaws in LLMs

### v1

@@ Abstract
Value model-guided search is effective in steering the generation but suffers from scaling flaws: Its superiority diminishes with larger sample sizes, underperforming non-search baselines. This limitation arises from reliability degradation in value models in unseen reasoning paths. To address this, we propose an uncertainty-aware search framework that includes two key components: (1) uncertainty-aware value models that incorporate uncertainty into predictions, and (2) an uncertainty-aware selection process using the proposed efficient Group Thompson Sampling algorithm. Experiments on GSM8K show that our method mitigates search scaling flaws, achieving 90.5% coverage at 16 samples compared to 85.8% for conventional value-guided search. This work establishes the first systematic integration of uncertainty quantification in LLM search paradigms.
Uncertainty-Aware Search and Value Models:
Mitigating Search Scaling Flaws in LLMs
@@
$\dagger$\$$\dagger$\$footnotetext: Corresponding to Yingru Li and Benyou Wang.
@@ 1 Introduction
(a) Scaling Sample Size
(b) Scaling Candidate Size
Figure 1: Comparison of scaling laws between OVM-guided search and UVM-guided search. Our proposed uncertainty-aware search framework alleviates search scaling flaws and coverages to higher coverage than the conventional VM-guided search.
Test-time scaling Brown et al. (2024); Snell et al. (2024); Wu et al. (2024) boosts performance significantly on multi-step mathematical reasoning tasks Cobbe et al. (2021); Hendrycks et al. (2021). Value Model (VM)-guided search Yu et al. (2024); Wan et al. (2024) efficiently solves more problems by steering the generation toward more effective reasoning paths.
However, a recent study Yu et al. (2025) identifies scaling flaws in conventional VM-guided search: it outperforms repeated sampling (i.e. non-search baseline) at small sample sizes but improves more slowly, leading to inferior performance at larger sample sizes. As shown in Table 1, VM-guided search surpasses repeated sampling at a sample size of 1 (75.4% v.s. 52.9%), but loses the superiority as the sample sizes increase to 8 and 16 (85.8% v.s. 90.8%).
Table 1: Comparison of coverage on GSM8K: Conventional VM-guided search faces search scaling flaws, underperforming the non-search baseline (repeated sampling). Our proposed uncertainty-aware search effectively enhances the effectiveness of the search scaling.
Sample Size 1 8 16
Repeated Sampling 52.9% ± 0.6% 84.7% ± 0.8% 90.8% ± 0.1%
VM-Guided Search 75.4% ± 0.6% 84.0% ± 0.4% 85.8% ± 0.3%
Uncertainty-Aware Search (Ours) 67.8% ± 0.4% 87.5% ± 0.6% 90.5% ± 0.1%
According to Yu et al. (2025), this issue arises from VM failures during the evaluation and selection of candidates in the search process. When VMs underestimate the values of promising candidates and misidentify them, the selection fails, ultimately leading to search failures.
@@
To mitigate scaling flaws, it is crucial to address unreliable VM predictions at the selection stage. Since VMs rely heavily on training data, they are more likely to provide unreliable predictions when encountering unseen data during inference. This can be addressed by incorporating uncertainty to reflect the reliability of VM evaluations, particularly for candidates underrepresented in the training data. Intuitively, candidates less frequently seen during training have higher uncertainty, indicating lower VM reliability and greater risk of selection failures.
To capture uncertainty and develop Uncertainty-Aware Value Models (UVMs), we employ Ensemble++ architecture Li et al. (2024a, b) to model value distributions, which encapsulate the inherent uncertainty, with a more dispersed distribution indicating higher uncertainty. The trained UVM allows us to evaluate candidates while accounting for uncertainty by sampling from the value distribution.
@@
Leveraging the accessibility to uncertainty-aware value distributions, we develop an efficient algorithm for uncertainty-aware candidate selection. We propose the Group Thompson Sampling algorithm, an innovative extension of Thompson Sampling Thompson (1933); Russo et al. (2018). This algorithm selects candidates based on their probability of being the optimal choice within the candidate set, i.e. top-1 probability. This method is effective and remains efficient, without the need for explicit top-1 probability estimation. As shown in Table 1 and Figure 1, our approach outperforms the conventional VM in search scaling effectiveness.
In summary, our contributions are as follows: (1) We introduce an uncertainty-aware search framework that incorporates uncertainty quantification into the selection stages during the search process, enhancing the search scaling effectiveness (2) We develop uncertainty-aware value models, equipping VMs with uncertainty quantification (3) We propose the Group Thompson Sampling algorithm, enabling efficient uncertainty-aware candidate selection.
@@ 2 Background
Scaling flaws
Scaling sample size $b$ or candidate size $K$ are expected to handle more problems. However, Yu et al. (2025) observed that search suffers from scaling flaws: (1) the performance increases at a slower rate than the non-search baseline as $b$ grows, ultimately converging to a much lower point (2) the performance remains unimproved and even degrades as $K$ scale.
These issues stem from the failures in the selection stage, where promising paths among candidates are not identified and are improperly pruned. These arise due to the use of imperfect value models for candidate evaluation Yu et al. (2025).
OVM and VM failures
The Outcome-supervised Value Model Yu et al. (2024) trains a value model to evaluate each candidate by predicting the probability of it reaching a correct answer. Then, it selects candidates with the highest predicted values from the set $\mathbb{S}^{(1:t)}$.
However, imperfect value models may misidentify and incorrectly rank promising paths, which leads to selection failures and search scaling flaws Yu et al. (2025).
Figure 2: Illustration of the UVM structure, value learning process, and its relationship to OVM: (i) This figure shows how UVM extends OVM by adding uncertainty with minimal additional parameters. The blue branch represents OVM, which computes a mean value for a sequence. The orange branch introduces the uncertainty term in UVM, calculated using parameters $\mathbf{W}$ and $\mathbf{W}_{0}$. This uncertainty term varies with the input index $\boldsymbol{\zeta}$, leading to diverse posterior value samples. The process of using UVM is simple: UVM derives a mean value like OVM, but also samples from a fixed distribution and adds the uncertainty term. (ii)-(iii) For training, UVM uses the same training set as OVM, but samples 2m posterior values $[v_{1},\dots,v_{2m}]$ using a discrete coordinate distribution $[\boldsymbol{e}_{1},\dots,\boldsymbol{e}_{2m}]$, rather than estimating a single value. The model is trained by averaging the MSE over these posterior samples.
@@ 3.1 Uncertainty-Aware Value Model
Ensemble++ Li et al. (2024a, b) is an ensemble-based approach that captures data uncertainty by modelling the posterior distribution. When test data resembles sufficiently seen data, the posterior distribution is concentrated; otherwise, it is more dispersed. This approach is simple to implement, requiring only a learnable linear transformation of the existing representation $\mathbf{x}$. It learns to map a predefined distribution $p_{\boldsymbol{\zeta}}$, like a Gaussian, to the target posterior distribution.
@@
We borrow Ensemble++ Li et al. (2024a, b) to model uncertainty-aware values as illustrated in Figure 2 (i). There are three processes involved in UVM: (1) representation encoding (2) index sampling (3) index mapping
Representation encoding: The last hidden states from a LLM backbone (parameterized by $\boldsymbol{\theta}$) serve as the representation $\mathbf{x}$. Specifically, for a given question $q$ and a partial path $S^{(1:t)}$, the representation $\mathbf{x}$ is obtained as:
$\mathbf{x}=\operatorname{LLM}(q,S^{(1:t)};\boldsymbol{\theta})$ (1)
Index sampling: This process samples an index from the predefined distribution, i.e. $\boldsymbol{\zeta}\sim p_{\boldsymbol{\zeta}}$
Index mapping: The index is mapped to the posterior value by summing a mean value term and an uncertainty term:  [^1] For simplicity, we present the main expression and omit the hyperparameters involved in the practical implementation. Details on the practical implementation can be found in A.2.
$v=\underbrace{\mathbf{x}\mathbf{b}}_{\text{mean estimator}}+\underbrace{\mathbf{x}(\mathbf{W}+\mathbf{W}_{0})\boldsymbol{\zeta}^{T}}_{\text{uncertainty estimator}}$ (2)
Through these processes, a trained UVM maps the predefined distribution $p_{\boldsymbol{\zeta}}$ to the posterior value distribution $p(v|q,S^{(1:t)})$.
@@
Here, $\mathbf{W}^{d\times m}$ and $\mathbf{b}^{d\times 1}$ are learnable parameters, while $\mathbf{W}_{0}^{d\times m}$ are frozen parameters that are randomly initialized. $d$ represents the dimension of the backbone’s hidden states $\mathbf{x}^{1\times d}$. $m$ is a hyperparameter for the dimension of the index $\boldsymbol{\zeta}^{1\times m}$. Notably, UVM is simple and straightforward to implement on top of OVM, requiring only an additional linear transformation.
@@
UVM can be regarded as a last-layer ensemble of $m$ components, controlled by the index vector $\boldsymbol{\zeta}^{1\times m}$: (1) When $\boldsymbol{\zeta}^{1\times m}$ is a zero vector $[0,\dots,0]$, UVM retains only the deterministic branch, degenerating to OVM (2) When $\boldsymbol{\zeta}^{1\times m}$ is a one-shot index vector, with a 1 in the $i$-th position and 0s elsewhere, it queries the $i$-th component for the prediction (3) When $\boldsymbol{\zeta}^{1\times m}$ is a non-one-shot vector, it is equivalent to use a linear combination of the $m$ components.
@@ 3.2 Accessing Values
To derive a posterior value, we (1) sample an index $\boldsymbol{\zeta}\sim p_{\boldsymbol{\zeta}}$ and (2) map it to the value sample $v$ using Equation 1-2. Notably, deriving multiple posterior values only requires one-time representation encoding, i.e. forward pass through the LLM backbone (Equation 1), followed by multiple posterior value mappings (Equation 2) with repeated index sampling and mapping.
@@
The mean is obtained by using a zero vector as the index $\boldsymbol{\zeta}$ and mapping it. The standard deviation is estimated by sampling multiple posterior values, which involves repeated index sampling and mapping.
@@ 3.3 Uncertainty-Aware Value Learning
UVM is learning a posterior value distribution, which complicates its learning (Figure 2 (iii)). Following Ensemble++ Li et al. (2024a, b), a discrete coordinate distribution is used in training. This distribution consists of $2m$ one-hot index vectors $[\boldsymbol{e}_{1},\dots,\boldsymbol{e}_{m},\boldsymbol{e}_{m+1},\dots,\boldsymbol{e}_{2m}]$. For each $i=1,\dots,m$, the $i$-position of $\boldsymbol{e}_{i}$ is 1, while of $\boldsymbol{e}_{m+i}$ is -1, with 0s elsewhere. The training loss for each $(q,S^{(1:T)},y)$ tuple is
$L^{\operatorname{UVM}}(q,S^{(1:T)},y)=\sum_{t=1}^{T}\frac{1}{2m}\sum_{i=1}^{2m}(\operatorname{UVM}(q,S^{(1:t)},\boldsymbol{e}_{i})-y)^{2}$ (4)
Notably, the gradients of the term $\mathbf{W}$ are stopped to prevent them from propagating to the backbone parameters $\boldsymbol{\theta}$.
@@ 4 Uncertainty-Aware Selection
Following Ensemble++ Li et al. (2024a, b), a $m$-dimensional continuous Gaussian distribution is used as the index distribution $p_{\boldsymbol{\zeta}}$ to access the expressive posterior value distribution during inference, as described in Section 3.2. Intuitively, we are training $m$ individual components during value learning, and then combine them linearly to make predictions during inference.
@@
We use top-1 probability for candidate selection instead of the well-known Upper Confidence Bound (UCB)  [^2] UCB scores each candidate using the sum of an exploitation term and an exploration term, i.e. the sum of mean values and standard variation in our case. for two main reasons:
@@
Group Thompson Sampling is an extension of the Thompson sampling algorithm, designed to efficiently select a group of candidates based on their top-1 probability. Thompson sampling Thompson (1933); Russo et al. (2018) is a straightforward method for selecting a single candidate, by performing one posterior sampling and choosing the one with the highest sampled value
@@
1:Input: Question $q$, candidates $\{S_{1},\dots,S_{K}\}$, sample size $b$
@@ 5 Experiment Results
This section outlines our experiment settings and presents the results of overall performance and ablation studies.
@@ 5.1 Experimental Settings
Benchmarks and models
We conduct experiments on GSM8K Cobbe et al. (2021) using Mistral 7B Jiang et al. (2023). We use the official training split and test split for all the experiments.
Baselines
We compare our method, UVM-guided search, with two baselines: (1) repeated sampling, which directly samples multiple solution paths without any search mechanism, and (2) conventional OVM-guided search, which does not incorporate uncertainty.
@@
Following the study of search scaling flaws Yu et al. (2025), we evaluate our method using coverage – the fraction of problems whose correct solutions for which the correct solution is covered by the generated paths, i.e., at least one sampled path is correct. This metric is also referred to as pass@k, where $k$ denotes the number of generated paths.  [^3] In our paper, $K$ represents the candidate size during the intermediate search process, rather than the number of final produced paths. To avoid confusion, we use the term “coverage”.
Scaling beam search
Yu et al. (2025) observes scaling flaws related to both sample sizes and candidate sizes. We also investigate the effectiveness of our methods on with respect to these two factors:
Sample size: This refers to the number of complete solution paths generated by the algorithm. For beam search, it corresponds to the beam size $b$, while for repeated sampling, it is the number of attempts. We consider sample sizes of 1, 2, 4, 8, 16, and 32. For beam search experiments, we fix the number of generated paths per beam $K/b$ at 8.
Candidate size: This represents the number of candidates considered during the search process. In these experiments, we fix the sample size at 16, where OVM-guided search underperforms repeated sampling. We conduct experiments with candidate sizes of 32, 64, 128, and 256.
Each experiment is repeated three times, and we report the average coverage along with its standard deviation.

### v2

# Robust Search with Uncertainty-Aware Value Models for Language Model Reasoning

@@ Abstract
Value model guided search is effective in steering LLM generation but suffers from a lack of robustness. This is due to verifier failure: imperfect VMs mistakenly prune valid reasoning paths, especially when encountering unseen reasoning paths generated during search. To address this, we propose an uncertainty-aware framework with two key components: (1) Uncertainty-Aware Value Models (UVMs), which replace single-point value estimates with value distributions to quantify prediction reliability, and (2) Group Thompson Sampling, an efficient algorithm that selects candidates based on their probability of being optimal. Experiments on two In-Distribution (ID) settings (GSM8K, MATH) and three Out-Of-Distribution (OOD) settings (e.g., AIME25, Minerva Math) show our method significantly mitigates verifier failure and boosts solution coverage, especially on OOD problems. This work provides the first systematic integration of uncertainty quantification into LLM search paradigms, enhancing robustness. The code is released at https://github.com/FreedomIntelligence/UVM.
@@ 1 Introduction
Test-time scaling (Brown et al., 2024; Snell et al., 2024; Wu et al., 2024) significantly boosts performance on multi-step language model reasoning tasks (Cobbe et al., 2021; Hendrycks et al., 2021). Value Model (VM)-guided search (Yu et al., 2024; Wan et al., 2024) is a particularly effective strategy, efficiently steering generation toward promising reasoning paths.
However, the robustness of this approach is undermined by a critical limitation known as verifier failure (Yu et al., 2025). Imperfect VMs, when encountering reasoning paths not seen in their training data, often underestimate the value of promising candidates. This leads to selection failures, where these valid paths are incorrectly pruned, ultimately causing the search to fail.
The root of this problem is that conventional VMs provide a single point estimate without conveying their confidence. A VM may be highly certain or merely guessing, but the search algorithm cannot tell the difference. Our core insight is that for a robust search, the algorithm must know when its value model is unreliable.
@@
To mitigate such failures, we introduce Uncertainty-Aware Value Models (UVMs). Instead of a single value, UVMs model a distribution over a path’s value. We employ the efficient Ensemble++ architecture (Li et al., 2024a; b) to achieve this. A dispersed distribution signals high uncertainty, warning the search algorithm not to prematurely discard a candidate based on an unreliable low-value estimate.
@@
Having access to value distributions requires a new selection strategy. We propose Group Thompson Sampling, an innovative and efficient extension of Thompson Sampling (Thompson, 1933; Russo et al., 2018). It selects candidates by sampling based on their top-1 probability (the probability of being the best choice in the set) without needing to compute this probability explicitly. This provides a principled mechanism to balance high-value and high-uncertainty candidates.
In summary, our contributions are as follows: (1) We address verifier failure by proposing a systematic uncertainty-aware framework for robust search. (2) We develop UVMs that efficiently quantify prediction uncertainty for language model reasoning paths. (3) We propose the Group Thompson Sampling algorithm for efficient and effective uncertainty-aware selection.
@@ 2 Background
OVM, VM failures, and selection failures
The Outcome-supervised Value Model (Yu et al., 2024) trains a value model to evaluate each candidate by predicting the probability of it reaching a correct answer. Then, it selects candidates with the highest predicted values from the set $\mathbb{S}^{(1:t)}$. However, the imperfect value model may misidentify and incorrectly rank promising paths, improperly pruning these promising paths, which leads to selection failures (Yu et al., 2025).
@@ 3.1 Uncertainty-Aware Value Model
Ensemble++ (Li et al., 2024a; b) is an ensemble-based approach that captures data uncertainty by modelling the posterior distribution. When test data resembles sufficiently seen data, the posterior distribution is concentrated; otherwise, it is more dispersed. This approach is simple to implement, requiring only a learnable linear transformation of the existing representation $\mathbf{x}$. It learns to map a predefined distribution $p_{\bm{\zeta}}$, like a Gaussian, to the target posterior distribution.
@@
We borrow Ensemble++ (Li et al., 2024a; b) to model uncertainty-aware values as illustrated in Figure 1 (i). There are three processes involved in UVM: (1) representation encoding (2) index sampling (3) index mapping
Representation encoding: The last hidden states from a LLM backbone (parameterized by $\bm{\theta}$) serve as the representation $\mathbf{x}$. Specifically, for a given question $q$ and a partial path $S^{(1:t)}$, the representation $\mathbf{x}$ is obtained as:
$\mathbf{x}=\operatorname{LLM}(q,S^{(1:t)};\bm{\theta})$ (1)
Index sampling: This process samples an index from the predefined distribution, i.e. $\bm{\zeta}\sim p_{\bm{\zeta}}$
Index mapping: The index is mapped to the posterior value by summing a mean value term and an uncertainty term: $\dagger$$\dagger\daggerFor$ simplicity, we present the main expression and omit the hyperparameters involved in the practical implementation. Details on the practical implementation can be found in A.2.
$v=\underbrace{\mathbf{x}\mathbf{b}}_{\text{mean estimator}}+\underbrace{\mathbf{x}(\mathbf{W}+\mathbf{W}_{0})\bm{\zeta}^{T}}_{\text{uncertainty estimator}}$ (2)
Through these processes, a trained UVM maps the predefined distribution $p_{\bm{\zeta}}$ to the posterior value distribution $p(v|q,S^{(1:t)})$.
@@
Here, $\mathbf{W}^{d\times m}$ and $\mathbf{b}^{d\times 1}$ are learnable parameters, while $\mathbf{W}_{0}^{d\times m}$ are frozen parameters that are randomly initialized. $d$ represents the dimension of the backbone’s hidden states $\mathbf{x}^{1\times d}$. $m$ is a hyperparameter for the dimension of the index $\bm{\zeta}^{1\times m}$. Notably, UVM is simple and straightforward to implement on top of OVM, requiring only an additional linear transformation.
@@
UVM can be regarded as a last-layer ensemble of $m$ components, controlled by the index vector $\bm{\zeta}^{1\times m}$: (1) When $\bm{\zeta}^{1\times m}$ is a zero vector $0,\dots,0]$, UVM retains only the deterministic branch, degenerating to OVM (2) When $\bm{\zeta}^{1\times m}$ is a one-shot index vector, with a 1 in the $i$-th position and 0s elsewhere, it queries the $i$-th component for the prediction (3) When $\bm{\zeta}^{1\times m}$ is a non-one-shot vector, it is equivalent to use a linear combination of the $m$ components.
Figure 1: Illustration of the UVM structure, value learning process, and its relationship to OVM: (i) This figure shows how UVM extends OVM by adding uncertainty with minimal additional parameters. The blue branch represents OVM, which computes a mean value for a sequence. The orange branch introduces the uncertainty term in UVM, calculated using parameters $\mathbf{W}$ and $\mathbf{W}_{0}$. This uncertainty term varies with the input index $\bm{\zeta}$, leading to diverse posterior value samples. The process of using UVM is simple: UVM derives a mean value like OVM, but also samples from a fixed distribution and adds the uncertainty term. (ii)-(iii) For training, UVM uses the same training set as OVM, but samples 2m posterior values $v_{1},\dots,v_{2m}]$ using a discrete coordinate distribution $\bm{e}_{1},\dots,\bm{e}_{2m}]$, rather than estimating a single value. The model is trained by averaging the MSE over these posterior samples.
@@ 3.2 Accessing Values
To derive a posterior value, we (1) sample an index $\bm{\zeta}\sim p_{\bm{\zeta}}$ and (2) map it to the value sample $v$ using Equation 1-2. Notably, deriving multiple posterior values only requires one-time representation encoding, i.e. forward pass through the LLM backbone (Equation 1), followed by multiple posterior value mappings (Equation 2) with repeated index sampling and mapping.
@@
The mean is obtained by using a zero vector as the index $\bm{\zeta}$ and mapping it. The standard deviation is estimated by sampling multiple posterior values, which involves repeated index sampling and mapping.
@@ 3.3 Uncertainty-Aware Value Learning
UVM is learning a posterior value distribution, which complicates its learning (Figure 1 (iii)). Following Ensemble++ (Li et al., 2024a; b), a discrete coordinate distribution is used in training. This distribution consists of $2m$ one-hot index vectors $\bm{e}_{1},\dots,\bm{e}_{m},\bm{e}_{m+1},\dots,\bm{e}_{2m}]$. For each $i=1,\dots,m$, the $i$-position of $\bm{e}_{i}$ is 1, while of $\bm{e}_{m+i}$ is -1, with 0s elsewhere. The training loss for each $(q,S^{(1:T)},y)$ tuple is
$L^{\operatorname{UVM}}(q,S^{(1:T)},y)=\sum_{t=1}^{T}\frac{1}{2m}\sum_{i=1}^{2m}(\operatorname{UVM}(q,S^{(1:t)},\bm{e}_{i})-y)^{2}$ (4)
Notably, the gradients of the term $\mathbf{W}$ are stopped to prevent them from propagating to the backbone parameters $\bm{\theta}$.
@@ 4 Uncertainty-Aware Selection
Following Ensemble++ (Li et al., 2024a; b), a $m$-dimensional continuous Gaussian distribution is used as the index distribution $p_{\bm{\zeta}}$ to access the expressive posterior value distribution during inference, as described in Section 3.2. Intuitively, we are training $m$ individual components during value learning, and then combine them linearly to make predictions during inference.
@@
We use top-1 probability for candidate selection instead of the well-known Upper Confidence Bound (UCB) $\dagger$$\dagger\daggerUCB$ scores each candidate using the sum of an exploitation term and an exploration term, i.e. the sum of mean values and standard variation in our case. for two main reasons:
@@
Group Thompson Sampling is an extension of the Thompson sampling algorithm, designed to efficiently select a group of candidates based on their top-1 probability. Thompson sampling (Thompson, 1933; Russo et al., 2018) is a straightforward method for selecting a single candidate, by performing one posterior sampling and choosing the one with the highest sampled value
@@
1:Input: Question $q$, candidates $\{S_{1},\dots,S_{K}\}$, beam size $b$
@@ 5 Experiment Results
This section outlines our experiment settings and presents the results of overall performance.
@@ 5.1 Experimental Settings
Benchmarks
We compare our method, UVM-guided search, with the conventional OVM-guided search, which does not incorporate uncertainty. We consider both ID and OOD scenarios, including traditional OOD and Randomized Token Noise OOD (RTN-OOD) settings.
  • ID settings: We run experiments on GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). For training, we use the official training split of each dataset; for evaluation, we use the official GSM8K test set and MATH500 (Lightman et al., 2024).
[…]
  • RTN-OOD: We define an RTN‑OOD setting in which the VM training distribution consists of sequences whose tokens are sampled from the tokenizer vocabulary. Evaluation is conducted on MATH500, referred to as RTN-MATH. This enables a controlled comparison to ID MATH, focused on extreme training set shift, by varying only the training distribution.
Models
On GSM8K, we use Mistral 7B (Jiang et al., 2023) and Qwen2.5-Math 7B (Yang et al., 2024); on MATH, we use Qwen2.5-Math 7B and Qwen3 8B (Yang et al., 2025). For all OOD settings, we use Qwen2.5-Math 7B. $\dagger$$\dagger\daggerQwen$3‑8B is adjusted to produce traditional Chain-of-Thought (CoT) in this paper; observing no gains over Qwen2.5‑Math, we therefore use Qwen2.5‑Math for all OOD settings.
@@
We evaluate our method using two metrics: Coverage and precision. Coverage measures the fraction of problems for which the correct solution is covered by the generated paths, i.e., at least one sampled path is correct. This is also known as pass@k, where $k$ denotes the number of generated paths. $\dagger$$\dagger\daggerIn$ our paper, $K$ represents the candidate size during the intermediate search process, rather than the number of final produced paths. To avoid confusion, we use the term “coverage”. For simplicity, precision is quantified under the majority voting strategy (Wang et al., 2023) as the fraction of problems for which the majority answer among the generated paths is correct.
We conduct the main experiments under the largest inference setting in  Yu et al. (2025), using a beam size of 32 and a candidate size of 256. Each experiment is repeated three times, and we report the average metrics along with their standard deviation.